OpenStudy (anonymous):

Consider 3 trials, each having the same probability of successes. Let X denote the total number of successes in these trials. If E(X) = 1.8, what is: a. the largest possible value of P{X=3}? b. the smallest possible value of P{X=3}? In both cases, construct a probability scenario that results in P{X=3} having the stated value. Would really appreciate some help on this one guys...I'm told the answer to A is 0.6 but why?

OpenStudy (anonymous):

they each have the same probability right? \[E(x)=P(x=0)\times 0+P(x=1)\times 1+P(x=2)\times 2+P(x=3)\times 3=1.8\]

OpenStudy (anonymous):

if all the first ones are zero, then \[P(x=3)\times 3=1.8\] so \[P(x=3)=0.6\]

OpenStudy (anonymous):

See that's interesting because that emulates perfectly a lot of the solutions. My only question (stupidly) is why: P(X=0)*0 + P(X=1)*1 .... Is that because the value of that probability i.e. take P(X=1) for example, is 1?

OpenStudy (anonymous):

Wait yes my question is why are the first ones zero. How do you simply jump to P(x=3)*3 = 1.8 ?